### v2

# EVT: Efficient View Transformation for Multi-Modal 3D Object Detection

Abstract
Multi-modal sensor fusion in bird’s-eye-view (BEV) representation has become the leading approach in 3D object detection. However, existing methods often rely on depth estimators or transformer encoders for view transformation, incurring substantial computational overhead. Furthermore, the lack of precise geometric correspondence between 2D and 3D spaces leads to spatial and ray-directional misalignments, restricting the effectiveness of BEV representations. To address these challenges, we propose a novel 3D object detector via efficient view transformation (EVT), which leverages a well-structured BEV representation to enhance accuracy and efficiency. EVT focuses on two main areas. First, it employs Adaptive Sampling and Adaptive Projection (ASAP), using LiDAR guidance to generate 3D sampling points and adaptive kernels. The generated points and kernels are then used to facilitate the transformation of image features into BEV space and refine the BEV features. Second, EVT includes an improved transformer-based detection framework, which contains a group-wise query initialization method and an enhanced query update framework. It is designed to effectively utilize the obtained multi-modal BEV features within the transformer decoder. By leveraging the geometric properties of object queries, this framework significantly enhances detection performance, especially in a multi-layer transformer decoder structure. EVT achieves state-of-the-art performance on the nuScenes test set with real-time inference speed.
1 Introduction
In the field of 3D object detection, LiDAR-camera fusion is a crucial approach for enhancing perception systems [17, 6, 44, 34, 26, 21, 1, 7, 43, 39, 40, 3]. LiDAR sensors provide precise depth information, which is crucial for accurate object positioning. In contrast, camera sensors capture color and texture, offering rich semantic context. By combining the complementary strengths of these two sensors, robust and reliable 3D object detection can be achieved even in complex environments. This capability is critical in autonomous driving, where consistent and reliable perception is essential.
Figure 1: Performance comparison of EVT and other methods on the nuScenes validation set. The FPS of all methods is measured on a single Tesla A100 GPU using the official implementations, excluding voxelization time.
In multi-modal 3D object detection, the most common approach is to use an ego-centric bird’s-eye-view (BEV) representation space [17, 26, 21, 1, 3]. Early studies employ a 2D-to-3D projection method, which leverages the estimated depth distributions to transform perspective-view images into the BEV space [29, 12, 18, 26, 21]. Although this method utilizes spatial information via pixel-wise depth distributions, its heavy reliance on the performance of the depth estimator limits its overall robustness. Moreover, spatial misalignment frequently occurs when rasterizing the projected 3D points onto the BEV grids.
In contrast, a 3D-to-2D approach [20, 42, 3] has been introduced, where predefined 3D points are projected onto image planes, and features are sampled using deformable attention [49]. However, these predefined points struggle to effectively capture features from regions of interest (ROIs). In addition, the BEV features of the transformed image fail to fully resolve the misalignment in the ray direction.
To address these challenges, we propose a novel 3D object detector via efficient view transformation (EVT), which leverages a well-structured BEV representation to enhance accuracy and efficiency. EVT focuses on two main areas: (1) adaptive sampling and adaptive projection (ASAP) framework, which is a 3D-to-2D projection-based view transformation method, and (2) an improved transformer-based detection framework, which includes a group-wise query initialization method and an enhanced query update framework for the transformer decoder.
The ASAP framework produces high-quality BEV feature maps from images by leveraging LiDAR guidance, without requiring additional networks, such as depth estimators or transformer encoders. ASAP consists of two main modules. The adaptive sampling (AS) module generates 3D points to transform features from ROIs in images into BEV space. The adaptive projection (AP) module then refines the BEV features obtained from the AS module using adaptive kernels, thereby enhancing feature representation and resolving ray-directional misalignment.
For the improved transformer-based detection framework, the query initialization method uses group-wise learnable parameters, obtaining significant performance improvements in a multi-layer decoder architecture compared to existing methods that initialize queries by sampling features [47, 1]. The enhanced query update framework further refines the deformable cross-attention [49] by integrating corner-aware sampling and position-embedded feature mixing. Corner-aware sampling improves feature sampling in deformable cross-attention by leveraging the geometric properties of object queries. Subsequently, position-embedded feature mixing aggregates the sampled features by explicitly embedding their spatial locations to provide a more detailed representation.
The proposed EVT presents several advantages over existing methods. First, multi-modal fusion in BEV space, without depth estimators or transformer encoders for view transformation, is efficient in terms of computational time and memory usage. Second, the improved transformer-based detection framework that utilizes the characteristics of class groups and geometric properties of objects offers a simple yet powerful solution that can be easily integrated with other transformer-based approaches. As shown in Fig. 1, the proposed EVT achieves 73.7% NDS and 8.3 FPS with ResNet-50 [10], 74.6% NDS and 4.9 FPS with V2-99 [15], and 71.7% NDS and 12.1 FPS with the LiDAR-only model EVT-L on the nuScenes validation set [2], outperforming other methods in both accuracy and inference time. On the nuScenes test set, EVT achieves 75.3% NDS, and 72.5% mAP without model ensemble or test-time augmentation, outperforming previous state-of-the-art methods.
In summary, the main contributions are as follows:
  • The proposed EVT is a novel 3D object detector that leverages a well-structured BEV representation to enhance accuracy and efficiency.
  • The proposed ASAP framework leverages LiDAR guidance to generate high-quality BEV feature maps without substantial computational overhead.
  • The improved transformer-based detection framework enhances overall detection performance and can easily be extended to other transformer-based detection methods.
  • EVT achieves state-of-the-art performance of 75.3% NDS and 72.6% mAP on nuScenes test set.
2 Related Work
Figure 2: Overall architecture of EVT. Each backbone extracts either image features or LiDAR features. The proposed ASAP module fuses these two features in BEV space. The group-wise query initialization stage generates queries from the fused feature map. In the transformer decoder, corner-aware sampling leverages the geometric properties of the object queries to sample multi-modal features, and position-embedded feature mixing decodes the sampled features to update the queries. These queries then predict the 3D bounding boxes.
2.1 Multi-modal Fusion in Transformer
Inspired by the pioneering work of DETR [5], numerous transformer-based detection methods [38, 24, 25, 28, 37, 13, 23, 6, 17, 43, 40] have been proposed. Meanwhile, in multi-modal 3D object detection, extracting features from multiple sensors is still a key challenge. Capturing the complementary relationships between the data from different sensors leads to slow convergence and high computational complexity. To address this issue, FUTR3D [6] proposes a modality-agnostic feature sampler to aggregate features from different sensors using deformable cross attention [49]. DeepInteraction [43] preserves modality-specific information by using a modality interaction strategy. CMT [40] constructs input tokens for the transformer decoder by adding 3D position embeddings to each sensor’s data and then concatenating them. Despite recent advancements, these methods adopt iterative feature fusion in multiple layers, resulting in substantial computational overhead, while the lack of camera parameter utilization limits the effective use of geometric information.
2.2 Multi-modal Fusion in BEV Space
Object detection using a bird’s-eye view (BEV) representation [30, 29, 12, 11, 18, 17, 26, 21, 19, 20, 42, 3] focuses on view transformation of 2D image features into BEV space. For view transformation, 2D-to-3D projection methods map multi-view image features into BEV space by predicting pixel-wise depth distributions [29, 12, 11, 18, 17, 26, 21]. These methods have been extensively studied due to their ability to incorporate spatial priors through depth estimation. BEVFusion [26] fuses LiDAR and camera features in BEV space to provide complementary and robust results. In contrast, 3D-to-2D projection methods project 3D points onto image planes for view transformation without leveraging depth information [19, 20, 42]. BEVFormer [20] projects predefined BEV queries and refines the query features using deformable cross-attention. BEVFusion4D [3] enhances BEV queries by concatenating LiDAR and image BEV features.
However, reliance on inaccurate depth estimation or the use of transformer encoders without fully leveraging geometric properties presents significant challenges in accurately capturing the correlation between 2D and 3D spaces. Furthermore, using additional networks results in substantial computational overhead.
2.3 Query for Transformer-based Detection
Query initialization is crucial in model design for enhancing the performance of transformer-based object detection. Typically, query initialization involves sampling queries based on a heatmap to represent potential object centers. CenterFormer [47] identifies high-scoring keypoints in the heatmap and samples features from the corresponding positions in the input data to initialize queries. TransFusion [1] further incorporates category embeddings into the sampled features for query initialization and fuses these queries with image features in transformer layers. However, this sampling strategy introduces ambiguity by capturing only parts of an object or including multiple object instances, as highlighted by DINO [45].
Furthermore, effectively updating queries within the transformer decoder is essential for enhancing the accuracy of object detection. DETR3D [38] projects 3D learnable queries onto image planes to directly sample the corresponding features. BEVFormer [20] assigns a set of predefined anchor heights to each query on the BEV plane to create 3D reference points and then predicts the 2D sampling offsets on the image planes after projection. SparseBEV [23] generates 3D sampling offsets to formulate 3D sampling points by applying bounding box size and rotation to each query, and compensates sampling points based on the object and ego motion across multiple frames. Although existing methods have made significant progress, there remains considerable potential to further leverage geometric properties for 3D object detection.
3 Methodology
The proposed EVT is a novel method for 3D object detection with effective multi-modal fusion and improved transformer-based detection framework. The overall pipeline is shown in Fig. 2. First, $N_{\text{s}}$-scale perspective-view image features $\{\mathbf{PV}_{j}\mid j\in\{1,2,\dots,N_{\text{s}}\}\}$ and bird’s-eye-view (BEV) LiDAR features $\mathbf{BEV}_{\text{lidar}}\in\mathbb{R}^{C\times H\times W}$ are extracted from individual backbones, where $C$ is the feature dimension, and $H\times W$ is the size of the BEV feature map. Next, the proposed adaptive sampling and adaptive projection (ASAP) module fuses the multi-modal features in BEV space (Sec. 3.1). Group-wise queries are then initialized in BEV space using the heatmap head (Sec. 3.2), and updated in the transformer decoder based on the geometric properties of the 3D bounding box (Sec. 3.3).
3.1 Adaptive Sampling and Adaptive Projection
The proposed ASAP module effectively leverages LiDAR guidance to transform multi-scale perspective-view image features into BEV space. The structure of this module is shown in Fig. 3.
Adaptive Sampling
The adaptive sampling (AS) module samples multi-scale image features into BEV space by predicting the height of each sampling point on the BEV plane and assigning weights to the predicted heights and image scales.
First, a set of heights for each grid cell is generated from LiDAR features:
$\{Z_{i}\}_{i=1}^{N_{\text{h}}}={\rm Conv}(\mathbf{BEV}_{\text{lidar}})(u,v),$ (1)
where $(u,v)$ denotes the coordinates of a grid cell in BEV space, and $N_{\text{h}}$ denotes the number of generated heights. Consequently, the 3D adaptive sampling points $P=\{(X,Y,Z_{i})\}_{i=1}^{N_{\text{h}}}$ are defined using real-world coordinates $(X,Y)$ corresponding to the grid cell at $(u,v)$ and the set of heights $\{Z_{i}\}$.
Next, the generated points $P$ are projected onto the image plane. Each projected point samples the multi-scale features $\{\mathbf{PV}_{j}\}_{j=1}^{N_{\text{s}}}$ across $N_{\text{s}}$ different image scales with downsampled strides $\{S_{j}\}_{j=1}^{N_{\text{s}}}$. Consequently, $N_{\text{h}}\times N_{\text{s}}$ projected points $(x^{j}_{i},y^{j}_{i})$ and sampled features $f^{j}_{i}$ are obtained from the projection:
$(x^{j}_{i},y^{j}_{i})=\frac{{\rm Proj}(X,Y,Z_{i})}{S_{j}}$ (2)
$f^{j}_{i}={\rm B}(\mathbf{PV}_{j},(x^{j}_{i},y^{j}_{i}))\in\mathbb{R}^{C},$ (3)
where ${\rm Proj}(\cdot)$ denotes the projection of 3D points onto the image plane, and ${\rm B}(\cdot)$ denotes bilinear interpolation.
To aggregate all the $N_{\text{s}}\times N_{\text{h}}$ sampled features $\{f^{j}_{i}\}$, adaptive sampling weights $W_{\text{as}}\in\mathbb{R}^{N_{\text{s}}\times N_{\text{h}}}$ are obtained from $\mathbf{BEV}_{\text{lidar}}$. $W_{\text{as}}$ determines the importance of the heights and image feature scales for each grid cell. The multi-scale image features in BEV space are obtained as follows:
$W_{\text{as}}=\sigma({\rm Conv}(\mathbf{BEV}_{\text{lidar}})(u,v))$ (4)
$\mathbf{BEV}_{\text{as}}(u,v)=\sum_{j=1}^{N_{\text{s}}}\sum_{i=1}^{N_{\text{h}}}W_{\text{as}}(j,i)\cdot f^{j}_{i},$ (5)
where $\sigma(\cdot)$ denotes the softmax function applied to all $N_{\text{s}}\times N_{\text{h}}$ elements. $\mathbf{BEV}_{\text{as}}$ denotes the image BEV feature map, representing the features of the ROIs from the images in BEV space by leveraging LiDAR guidance.
Figure 3: Overview of AS and AP. In AS, the LiDAR BEV feature map generates 3D sampling points and their corresponding weights. The sampling points are projected onto the image plane, where features are sampled and combined by weighted pooling in each BEV grid cell. In AP, the image BEV feature map produced by AS is refined channel-wise using the adaptive kernels generated from the LiDAR features.
Adaptive Projection
The adaptive projection (AP) module refines the BEV feature $\mathbf{BEV}_{\text{as}}$ produced by the AS module using adaptive kernels for each grid cell. The overall process is represented by the following equations:
$K_{\text{ap}}={\rm Conv}(\mathbf{BEV}_{\text{lidar}})(u,v)\in\mathbb{R}^{C\times C}$ (6)
$\mathbf{BEV}_{\text{camera}}(u,v)=\mathbf{BEV}_{\text{as}}(u,v)\times K_{\text{ap}}\in\mathbb{R}^{C}.$ (7)
First, an adaptive kernel $K_{\text{ap}}\in\mathbb{R}^{C\times C}$ is generated from $\mathbf{BEV}_{\text{lidar}}$. This kernel is generated for each BEV grid cell to enable detailed feature representation. Subsequently, $\mathbf{BEV}_{\text{as}}$ is refined by applying a channel-wise linear projection using $K_{\text{ap}}$ to obtain the image BEV feature map $\mathbf{BEV}_{\text{camera}}$.
This LiDAR-guided refinement improves feature representation and addresses the ray-directional misalignment inherent in 3D-to-2D projection methods, which suffer from occlusions or empty 3D spaces.
Multi-modal Fusion in BEV Space
The multi-modal BEV feature map $\mathbf{BEV}_{\text{fuse}}\in\mathbb{R}^{C\times H\times W}$ is obtained by channel-wise concatenation of the image and LiDAR BEV feature maps, followed by a convolution operation:
$\mathbf{BEV}_{\text{fuse}}={\rm Conv}({\rm Concat}(\mathbf{BEV}_{\text{camera}},\mathbf{BEV}_{\text{lidar}})),$ (8)
where ${\rm Concat}(\cdot)$ denotes the channel-wise concatenation.
Method Modality NDS (val) mAP (val) NDS (test) mAP (test)
UVTR-L [17] L 67.7 60.9 69.7 63.9
TransFusion-L [1] L 70.1 65.1 70.2 65.5
FocalFormer3D-L [7] L - - 72.6 68.7
CMT-L [40] L 68.6 62.4 70.1 65.3
EVT-L (Ours) L 71.7 66.4 72.1 67.7
MVP [44] LC 70.8 67.1 70.5 66.4
UVTR [17] LC 70.2 65.4 71.1 67.1
AutoAlignV2 [8] LC 71.2 67.1 72.4 68.4
TransFusion [1] LC 71.3 67.5 71.7 68.9
DeepInteraction [43] LC 72.6 69.9 73.4 70.8
BEVFusion [26] LC 71.4 68.5 72.9 70.2
Objectfusion [4] LC 72.3 69.8 73.3 71.0
FocalFormer3D [7] LC 71.1 66.5 73.9 71.6
CMT [40] LC 72.9 70.3 74.1 72.0
BEVFusion4D-S [3] LC 72.9 70.9 73.7 71.9
SparseFusion [39] LC 72.8 70.4 73.8 72.0
MSMDFusion [14] LC - - 74.0 71.5
UniTR [35] LC 73.3 70.5 74.5 70.9
EVT (Ours) LC 74.6 72.1 75.3 72.6
Table 1: Performance comparison on the nuScenes validation and test sets. The results are obtained without model ensemble or test-time augmentation. ‘L’ and ‘C’ denote LiDAR and Camera, respectively.
3.2 Group-wise Query Initialization
The proposed group-wise query initialization method effectively generates queries for transformer-based detection frameworks. First, group-wise heatmaps are predicted from the multi-modal BEV feature map $\mathbf{BEV}_{\text{fuse}}$. The groups are predefined, with each group containing similarly-sized classes. The predicted heatmaps have scores ranging from 0 to 1, representing the likelihood that each BEV pixel corresponds to the center of an object. The heatmap head is supervised by 2D Gaussian distributions centered at each object’s location. Next, the top-k keypoints are selected from each heatmap group. The positions of these keypoints are used as reference points for queries in BEV space.
This method initializes each query feature using group-wise learnable parameters. Queries within the same group share these parameters, which enables them to effectively capture the characteristics of group. This demonstrates better performance in the multi-layer transformer decoder than traditional query initialization approaches that sample features directly from the input feature map [47, 1]. Further details are provided in Sec. 4.4.
3.3 Enhanced Query Update Framework
For more effective query updates within a transformer decoder based on deformable attention [49], we propose the corner-aware sampling and position-embedded feature mixing methods.
Corner-aware Sampling.
The proposed corner-aware sampling method shifts the sampling reference from the center of the object to each corner and rotates sampling offsets according to the object’s heading. This approach addresses the limitations of existing center-based sampling methods, which generate sampling offsets around the object’s center, resulting in difficulties capturing the full spatial extent and shape of the objects.
First, the initial points $\{(\Delta x_{i}^{\prime},\Delta y_{i}^{\prime})\}$ are generated from the query feature $\mathbf{q}$ using a linear layer:
$\{(\Delta x_{i}^{\prime},\Delta y_{i}^{\prime})\mid i\in 0,1,...,N_{\text{p}}-1\}={\rm Linear}(\mathbf{q}),$ (9)
where $N_{\text{p}}$ denotes the number of sampling points.
Subsequently, the sampling offsets $\{(\Delta x_{i},\Delta y_{i})\}$ and sampling points $\{(x_{i},y_{i})\}$ are determined via a geometric transformation as follows:
$\left[\!\!\begin{array}[]{c}\Delta x_{i}\\[5.0pt] \Delta y_{i}\\ \end{array}\!\right]\!=\!\left[\!\!\begin{array}[]{cc}\cos\theta\!\!\!&-\!\sin\theta\!\!\\[5.0pt] \sin\theta\!\!\!&\cos\theta\!\!\!\!\\ \end{array}\!\!\right]\!\!\left[\!\!\begin{array}[]{c}{\rm I}_{j}\cdot\frac{l}{2}+\Delta x_{i}^{\prime}\\[5.0pt] {\rm I}_{j}^{\prime}\cdot\!\frac{w}{2}\!+\Delta y_{i}^{\prime}\\ \end{array}\!\!\right]\!$ (10)
$\left[\!\!\begin{array}[]{cc}x_{i}&y_{i}\!\end{array}\!\right]\!=\!\left[\!\!\begin{array}[]{cc}x_{c}&y_{c}\end{array}\!\!\right]\!+\!\left[\!\begin{array}[]{cc}\!\Delta x_{i}&\Delta y_{i}\!\!\\ \end{array}\!\right],$ (11)
where $(I_{j},I^{\prime}_{j})\,\in\{(1,1),(1,-1),(-1,1),(-1,-1)\}$. $j$ denotes the index of each corner and is the remainder of $i$ divided by four, the number of corners. $(x_{c},y_{c})$ denote the coordinates of the object’s center. $l$, $w$, and $\theta$ denote the length, width, and yaw of the predicted bounding box, respectively, obtained from the regression head of the previous transformer decoder layer. In the first decoder layer, where there is no preceding layer to provide bounding boxes, the equation is satisfied by setting $l$, $w$, and $\theta$ to zero.
The features are sampled by bilinear interpolation at each sampling point on the multi-modal fused features $\mathbf{BEV}_{\text{fuse}}$ as follows:
$g_{i}={\rm B}(\mathbf{BEV}_{\text{fuse}},(x_{i},y_{i}))\in\mathbb{R}^{C},$ (12)
where ${\rm B}(\cdot)$ denotes the bilinear interpolation. $g_{i}$ represents the sampled feature that captures important key tokens by leveraging the geometric properties of the query’s bounding box.
Position-embedded Feature Mixing.
Accurate locations of the sampled features are essential for providing a more detailed representation of the bounding boxes. In AdaMixer [9], the adaptive channel and spatial mixing are combined to leverage query guidance and effectively decode sampled features. However, this method is less effective for decoding the outputs of the corner-aware sampling, as shown in Tab. 3 (e), which involves complex geometric transformations of the sampling points. To address this, we propose position-embedded feature mixing, an enhancement based on AdaMixer that improves the decoding capability by incorporating embeddings for sampling offsets.
First, the sampling offsets are embedded as position vectors using sinusoidal position encoding [33], followed by a linear layer. Then, the position-embedded sampled features $G\in\mathbb{R}^{N_{\text{p}}\times C}$ are obtained by element-wise addition of the sampled feature $g_{i}$ and the position vector $e_{i}$:
$e_{i}$ $={\rm Linear}(\Phi_{pos}((x_{i},y_{i})))\in\mathbb{R}^{C}$ (13)
$G$ =$=\{(g_{i}+e_{i})\mid i\in{0,1,...,N_{\text{p}}-1}\},$ (14)
where $\Phi_{pos}(\cdot)$ denotes sinusoidal position encoding.
Subsequently, adaptive channel mixing is applied to $G$ using the dynamic weights $W_{c}$ generated from the query feature $\mathbf{q}$ to obtain the channel-mixed feature $G_{c}$:
$W_{c}$ $={\rm Linear}(\mathbf{q})\in\mathbb{R}^{C\times C}$ (15)
$G_{c}$ $={\rm ReLU}({\rm LN}(G\times W_{c})).$ (16)
Next, adaptive spatial mixing is applied to the spatial dimensions of $G_{c}$ using dynamic weights $W_{s}$ to obtain the spatial-mixed feature $G_{cs}$:
$W_{s}$ $={\rm Linear}(\mathbf{q})\in\mathbb{R}^{N_{\text{p}}\times N_{\text{p}}}$ (17)
$G_{cs}$ $={\rm ReLU}({\rm LN}(G_{c}^{T}\times W_{s})).$ (18)
Finally, the query feature is updated as follows:
$\mathbf{q^{\prime}}=\mathbf{q}+{\rm Linear}({\rm Flatten}(G_{cs})),$ (19)
where $\mathbf{q^{\prime}}$ represents the updated query feature that is enhanced for a more accurate feature representation by incorporating the positional information of each sampled point.
4 Experiments
Figure 4: Visualization of projected sampling points for each object. The top row shows the projections of predefined 3D points, and the bottom row shows the projections of the points generated by the AS module. Points of each object are denoted by different colors.
Figure 5: Comparison of BEV feature maps from different methods: (a) Vanilla, (b) AS, (c) AP, and (d) ASAP. Vanilla shows less informative and unaligned features. AS improves feature representation using adaptive sampling. AP corrects ray-directional misalignment. ASAP integrates both AS and AP, leading to the most refined and well-aligned feature representation.
4.1 Implementation Details
For the image backbone, either ResNet-50 [10] with a resolution of $704\times 256$ or V2-99 [15] with $1600\times 640$ is used with the feature pyramid network (FPN) [22]. For the LiDAR backbone, VoxelNet [46] is used with an ROI of $[-54.0m,54.0m]$ along $(X,Y)$ and $[-5.0m,3.0m]$ along $Z$, with a voxel size of $(0.075m,0.075m,0.2m)$. In ASAP, we use 4 points for each grid cell and the last two stages of each image backbone for view transformation. The size of the unified BEV feature map size is $180\times 180$. The object categories are divided into 6 groups, with 150 queries per group, resulting in a total of 900 queries. The transformer decoder has 6 layers, and the corner-aware sampling generates 16 sampling points per query. The feature dimension is set to 256.
Our model is trained on 8 RTX 3090 GPUs with a batch size of 16. The model is trained end-to-end for 10 epochs using CBGS [48], whereas GT sample augmentation [41] is applied for the first 9 epochs. The query denoising strategy [16] is also adopted. Gaussian Focal loss [36], Focal loss [31] and L1 loss are used for heatmap prediction, classification and regression, respectively. The AdamW [27] optimizer is adopted with a learning rate of $1\times 10^{-4}$ and a weight decay of $1\times 10^{-2}$ with a cyclical learning rate policy [32]. No model ensemble or test-time augmentation is applied during inference.
4.2 Dataset and Metric
[…]
Performance on this dataset is assessed using metrics like mean Average Precision (mAP), calculated over distance thresholds of 0.5, 1, 2, and 4m across all classes, and the nuScenes Detection Score (NDS), which offers a holistic evaluation by combining mAP with measures of translation, scale, orientation, velocity, and attribute errors.
4.3 Comparison with State-of-the-art Methods
As shown in Tab. 1, we compare EVT and its LiDAR-only model, EVT-L, with existing methods on the nuScenes validation and test sets. The multi-modal EVT achieves 75.3% NDS and 72.6% mAP, outperforming all existing state-of-the-art approaches on both the nuScenes validation and test sets. In particular, it surpasses recent methods, such as UniTR [35] by 0.8% NDS and 1.7% mAP, MSMDFusion [14] by 1.3% NDS and 1.1% mAP, and SparseFusion [39] by 1.5% NDS and 0.6% mAP. Furthermore, EVT shows a performance improvement of 3.2% NDS and 4.9% mAP compared with the LiDAR-only model EVT-L. In contrast, TransFusion [1] shows only a 1.5% NDS and 3.4% mAP improvement over its LiDAR-only model, TransFusion-L. This indicates that EVT effectively utilizes camera data through the proposed view transformation method.
We also compare EVT-L with the LiDAR-only versions of other multi-modal 3D object detectors. On the nuScenes validation set, EVT-L surpasses TransFusion [1] and CMT [40] by 1.6% and 3.1% NDS, respectively. EVT-L also demonstrates competitive performance on the test set. These results validate the effectiveness of the proposed query initialization and cross-attention mechanisms in enhancing 3D object detection.
LiDAR Camera AS AP NDS mAP
(a) 71.7 66.4
(b) 72.7 69.1
(c) 73.5 70.6
(d) 73.7 70.7
Table 2: Ablation study of the proposed VT method. The proposed ASAP shows a significant performance improvement compared to the vanilla VT method (b), which projects predefined 3D points.
Figure 6: Comparison of query initialization strategies on the nuScenes validation set. For the multi-layer transformer decoder, group-wise query initialization without feature sampling achieves the best performance.
Attention Reference scale rotate feature mixing position-embedded NDS mAP
(a) Standard Center 69.6 65.3
(b) Deformable Center 70.0 64.8
(c) 70.5 65.2
(d) Ours Corner 71.3 65.7
(e) 71.2 65.6
(f) 71.7 66.4
Table 3: Comparison of the cross-attention methods within the transformer decoder. The proposed corner-aware sampling and position-embedded feature mixing achieve significant performance improvements.
Initial Query Formulation NDS mAP
BEV Reference Point 71.7 66.4
Bbox from regression head 71.4 66.3
Bbox from learnable parameters 70.9 66.0
Table 4: Comparison of the initial query formulations for the first layer of the transformer decoder.
4.4 Ablation Studies
In this section, we describe the validation of each component of the proposed method on the nuScenes validation set. Unless otherwise specified, all experiments are conducted using the proposed LiDAR-only model EVT-L, trained for 10 epochs with CBGS [48] and the query denoising strategy [16].
Adaptive Sampling and Adaptive Projection
Tab. 2 displays the results of the ablation conducted on ASAP by retraining the entire model. In these experiments, ResNet [10] is used as the backbone network with a resolution of $704\times 256$. (a) shows the performance of EVT-L, the LiDAR-only model. In (b), the vanilla view transformation (VT) method, without LiDAR guidance, employs a 3D-to-2D projection of predefined 3D points for feature sampling.
In (c), AS achieves improvements of 0.8% NDS and 1.5% mAP compared to the vanilla VT method. The sampling points of the vanilla VT and AS are visualized in Fig. 4. The sampling points of AS are adaptively generated in the ROIs.
In (d), the entire ASAP achieves improvements of 1.0% NDS and 1.6% mAP compared to the vanilla VT method (b). The BEV feature maps of each component are visualized in Fig. 5. ASAP effectively transforms image features into BEV space and resolves ray-directional misalignment.
Group-wise Query Initialization
Different query initialization methods using the heatmap head are ablated in Fig. 6. All methods utilize a top-K keypoints selection approach to define queries, but differ in their specific implementations. (a) directly samples features from the input BEV feature map and combines them with group-wise embedding, which is similar to the approach commonly used in 3D object detectors [47, 1]. (b) initializes queries with group-wise embedding using global top-K selection. (c) is our proposed group-wise query initialization strategy, which initializes queries using top-K selection and embedding for each group.
In the results, the direct feature sampling approach in (a) is more effective in a single-layer transformer decoder, outperforming (c) by 0.6% NDS. However, in a multi-layer transformer decoder, this leads to performance degradation. Specifically, with 6 layers, (c) surpasses (a) by 1.0% NDS. Additionally, group-wise top-K (c) consistently outperforms global top-K (b) across all layers by approximately 0.1% NDS.
Enhanced Query Update Framework
We ablate the corner-aware sampling method, as demonstrated in Tab. 3 (a)-(d). (a) is the standard multi-head cross attention [33], and (b)-(c) are the standard deformable attention [49]. Specifically, in (c), the sampling offsets are scaled according to the bounding box size and rotated according to the heading. (d) is the proposed corner-aware sampling method, which samples features from the bounding box corners of each query and rotates them according to the heading. The corner-aware sampling method (d) achieves improvements of 0.8% NDS and 0.5% mAP compared to method (c).
The results of the ablation study on the feature mixing method are displayed in Tab. 3 (e)-(f). In (e), the standard feature mixing method of AdaMixer [9] does not yield any improvements over our method. In (f), the proposed position-embedded feature mixing demonstrates improvements of 0.4% NDS and 0.7% mAP compared to (d).
Initial Query Formulation for Transformer
As described in Sec. 3.3, the corner-aware sampling is applied starting from the second transformer decoder layer. We experimented on two different approaches to incorporate the corner-aware sampling into all transformer decoder layers. The first approach involves adding regression heads during the query initialization stage (Sec. 3.2) to predict the bounding boxes, whereas the second approach uses learnable bounding boxes from SparseBEV [23]. However, as shown in Tab. 4, neither of these approaches led to any noticeable performance improvement. Furthermore, the experiment using learnable bounding boxes results in performance degradation.
5 Conclusion
In this paper, we have proposed EVT, a novel transformer-based 3D object detector. The proposed view transformation method, ASAP, enables robust and efficient LiDAR-camera fusion by leveraging LiDAR guidance. In addition, the proposed group-wise query initialization method provides effective queries for a multi-layer transformer decoder. The corner-aware sampling and position-embedded feature mixing methods effectively utilize geometric properties to update queries, and these methods can easily be extended to other models. We expect EVT to provide valuable insights into multi-modal 3D object detection.